# FLrce: Resource-Efficient Federated Learning with Early-Stopping Strategy

### v2

@@ 4 Experiment
4.1 Experiment Setup
  • Extended MNIST (EMNIST) [38] is an image dataset consisting of 814,255 handwritten characters from 62 categories (10 digits and 52 upper/lower-case letters). Each sample is an black-and-white-based image with $28\times 28$ pixels.
  • CIFAR10 [39] is an image dataset consisting of 50,000 samples covering 10 classes of real-world objects. Each sample is an colorful image with $32\times 32$ pixels.
  • CIFAR100 [39] is an image dataset consisting of 50,000 samples covering 100 classes of real-world objects. Each sample is an colorful image with $32\times 32$ pixels.
  • Google Speech [34] is a voice dataset consisting of 101,012 samples covering 35 human spoken words. Each sample is a short audio of a word pronounced by a human.
Parameter settings. For EMNIST and Google Speech, we use a convolutional neural network (CNN) with two convolutional layers and one fully-connected layer [26]. For CIFAR10 and CIFAR100, we use a CNN with two convolutional layers and three fully-connected layers [27]. For all datasets, the number of global iterations is $T=100$ and the number of clients’ local training epoch is 5. The batch size is set to 128 for CIFAR10 and CIFAR100, and 16 for EMNIST and Google Speech. The learning rate $\eta$ is set to 2e-4 for EMNIST, 5e-4 for Google Speech, 0.1 for CIFAR10 and 0.5 for CIFAR100. The explore-exploit factor $\phi$ depends on round $t$, that is, the explore possibility of FLrce is 1.0 initially and decays by a coefficient of 0.98 every round.
System implementation. The experiment platform is built upon the Flower framework (1.4.0) [40] and Pytorch 2.0.0 [41]. We build a virtual network consisting of one server and 100 clients, with 10 clients participating in FL per round [25]. The server runs on a personal desktop computer, and selects clients through the application programming interface (API) of Flower. Clients run the local training program on NVIDIA Jetson Nano Developer Kits [42]. We allow the edge devices to switch identities between several clients to conduct large-scale experiments with a limited number of physical devices [26]. To create a training environment with heterogeneous client contributions, data are unevenly separated into clients following a Dirichlet distribution with a parameter of 0.1 [35].
Baselines. As mentioned before, existing efficient FL methods can be divided into three categories: message compression, accuracy relaxation and dropout. Therefore, we compare FLrce against the typical works in these three categories. Out of resource concerns, we do not include any works that simply address client heterogeneity without taking efficiency into account. The experiment baselines include:
  • Fedcom [16] let clients compress the parameter updates before uploading the updates to the server so as to reduce the communication costs.
Metrics. We comprehensively compare FLrce with the baselines in terms of accuracy, computation efficiency and communication efficiency. For accuracy, by the end of every global iteration, we test the global model on every local dataset and record the mean accuracy. For computation efficiency, we measure the overall energy consumption of the Jetson Nano Develop Kits during the local training process, and we combine the energy consumption with the final accuracy to derive the computation efficiency. For communication efficiency, we measure the bandwidth usage for transmitting the model between clients and the server, and we combine the bandwidth usage with the final accuracy to derive the communication efficiency.
@@ 4.1 Experiment Setup
4.2 Experiment Results
4.2.1 Accuracy and Learning efficiency
The experiment outcome shows that FLrce significantly outperforms the baselines. From Figure 10, we can see that the learning efficiency of FLrce is much higher than the baselines, as FLrce can achieve higher accuracy given the same rounds of training. Furthermore, with the help of the ES mechanism, FLrce is able to complete the global training task using only $40\sim 60$% of the maximum global iteration and still achieve an equivalent/higher final accuracy. Specifically, as shown by Table III, it takes only 44, 52, 63 and 47 rounds for FLrce to reach an accuracy of 76.1%, 58.6%, 53.8% and 36.6% on EMNIST, Google Speech, CIFAR10 and CIFAR 100 respectively. Comparatively, Fedcom uses 100 rounds to reach an accuracy of 78.2%, 55.8%, 42.9% and 32.3%. Fedprox uses 100 rounds to reach an accuracy of 65.9%, 47.1%, 39.5% and 29.9%. Dropout uses 100 rounds to reach an accuracy of 8.5%, 21.6%, 14.9% and 13.7%. Even with less training rounds, FLrce still achieves the highest accuracy on Google Speech, CIFAR10 and CIFAR100 and the second highest accuracy on EMNIST (second to Fedcom). On average, FLrce improves the accuracy by at least 3.9% compared with the baselines.
4.2.2 Computation efficiency
We use energy consumption as a measurement of the computation cost [10, 14]. During the local training process, We measure the energy consumption of the Jetson Nano Developer Kits using a plug-in power monitor [43], when the training is completed, we sum up all clients’ energy consumption to derive the computation cost of the entire system, the results are available in Figure 11. As shown by Figure 11, FLrce consumes the least energy on all datasets, demonstrating the strength of FLrce in reducing the computation costs. Furthermore, motivated by [14], we combine accuracy and energy consumption to explore the computation efficiency of the benchmarks. Given the final accuracy and energy consumption, the computation efficiency is given by Equation (8):
As defined by Equation 8, the computation efficiency is proportional to accuracy and the inverse of energy consumption. It represents how a FL method is capable of utilizing the given computation resource to train an accurate global model. For a FL method, the higher accuracy it achieves and the less energy it consumes, the more computation efficient it will be. Figure 12 shows the computation efficiency of the benchmarks, for better visualization, the results have been normalized. As shown by Figure 12, FLrce obtains the highest computation efficiency on all of the four datasets. On average, FLrce improves the relative computation efficiency by at least 47%. This result demonstrates FLrce’s ability to achieve remarkable performance even with limited computation resources.
@@ 4.2.2 Computation efficiency
EMNIST 76.1% 44 79.1% 83 N/A N/A
Google Speech 54.7% 36 58.6% 52 N/A N/A
CIFAR10 53.8% 63 54.2% 81 54.8% 85
CIFAR100 36.6% 47 35.4% 48 35.6% 69
TABLE IV: Round and accuracy corresponding to different values of the Early Stopping threshold $\psi$.
@@
Figure 15: Effect of the Early Stopping threshold on computation efficiency, data have been normalized for better visualization.
@@
Figure 16: Effect of the Early Stopping threshold on communication efficiency, data have been normalized for better visualization.
4.2.3 Communication efficiency
From Figure 13, we can see that FLrce has the least bandwidth usage on EMNIST, Google Speech and CIFAR100. For CIFAR10, FLrce has the second least bandwidth usage that is slightly higher than Dropout. Even though Dropout uses less bandwidth than FLrce on CIFAR10, the final accuracy of Dropout is much lower than FLrce (14.9% vs. 53.8%). In this case, Dropout does not achieve a high communication efficiency as it does not maintain accuracy. Figure 14 shows that FLrce outperforms the baselines in terms of communication efficiency. On average, FLrce improves the relative communication efficiency by at least 43%.
4.3 Effect of the Early Stopping Threshold
To study the effect of the ES threshold $\psi$, we run FLrce multiple times with different values of $\psi$. As defined by subsection 3.3, to make sure that FLrce properly stops at a point with an acceptable validation accuracy, the ES mechanism should only be activated when a relatively large degree of conflict occurs among the selected clients, that is, the value of $\psi$ should not be too small. Furthermore, $\psi$ should not be too large as well; otherwise the ES mechanism might not be activated in time and cannot reduce the resource consumption. Therefore, we tune $\psi$ with the values of 5.0, 5.5, 6.0, and the results are shown by Figure 15, Figure 16 and Table IV. Note that Figure 15 and Figure 16 look exactly the same. This is because the computation/communication cost of FLrce per round is expected to be the same, so the computation/communication efficiency is proportional to $\frac{\text{accuracy}}{\text{round}}$ (see Equations (8) and (9)). Therefore, after normalization, the relative values of the computation and communication efficiency both become $\frac{\text{accuracy}}{\text{round}}$. For better illustration, we plot Figures 15 and 16 separately even though they show the same values. As shown by Table IV, FLrce successfully stops on all Datasets with $\psi=$ 5.0 or 5.5. For EMNIST and Google Speech, FLrce fails to stop before the maximum global iteration with $\psi\geq$ 6. Therefore, to make sure the ES mechanism be successfully activated, $\psi$ should be selected between $5\sim 5$ 5.5. As shown by Figure 15 and Figure 16, the value of $\psi$ to achieve the highest computation/communication efficiency is 5.0 (i.e. $\frac{P}{2}$, half of the number of selected clients per round). Accordingly, we set $\psi$ to 5.0 for EMNIST, CIFAR10, CIFAR100. For Google Speech, we set $\psi$ to 5.5 as it achieves higher accuracy than $\psi=5.0$ while still achieving the highest efficiency compared with the baselines (see Figures 12 and 14). Based on the empirical results, we suggest that resource-constrained IoT systems use a relatively small $\psi$ ($\approx$50of the selected clients) to maximize the efficiency of utilizing of the limited resources. On the other hand, IoT systems with sufficient resource budget can use a relatively large threshold ($\approx$55$\sim$60of the selected clients) to achieve higher model accuracy. We aim to find the optimal value of $\psi$ via theoretical analysis in the future work.
@@ 4.3 Effect of the Early Stopping Threshold
4.4 Effect of the Early Stopping Mechanism
To investigate whether the ES mechanism will cause undesirable accuracy loss, we compare the final accuracy of FLrce and FLrce without early stopping (FLrce w/o ES), using the threshold $\psi$ specified in the previous subsection. The only difference between FLrce and FLrce w/o ES is that the latter ignores the ES criterion and keeps training until reaching the maximum global iteration, and that relationship-based client selection strategy still applies in both settings. The comparison between FLrce and FLrce w/o ES is shown in Figure 10, Figure 17 and Figure 18. Note that Figures 17 and 16 look identically for the same reason as stated in the previous subsection. From Figures 17 and 18, we can see that FLrce can significantly increase the efficiency of the resource utilization with marginal accuracy sacrifice. For EMNIST, Google Speech and CIFAR10, FLrce degrades the final accuracy by 3.6%, 5.2% and 2.4% respectively compared with FLrce w/o ES, while the efficiency of FLrce w/o ES is only about 4666similar-toabsent66$\sim$ 66∼ 66% of FLrce. Besides, compared with existing efficient FL frameworks, FLrce still maintains competitive performance even with the accuracy deduction, as shown by Figure 10 and Table III. In some cases, FLrce even improves the final accuracy, as shown by Figures 16(d) and 17(d), the accuracy of FLrce is 1.8% higher than FLrce w/o ES on CIFAR100. The possible reason is that FLrce prevents the over-fitting of FL by terminating the FL process before the global model converge to some local sub-optimal solution.
@@ 4.5 Discussion
4.5.1 FLrce minimizes the accuracy loss caused the trade-off between accuracy and efficiency
@@
For comparison, FLrce tackles this bottleneck from a novel perspective, while existing works reduce the resource consumption by reducing the computation/communication cost per round, FLrce reduces the overall resource consumption by reducing the total rounds of training. The advantage of this method is that FLrce does not have to make any trade-off (compress gradients, less epoch, etc.) in the training process, and can achieve a high accuracy faster given the same training rounds as the baselines. Based on this point, when FLrce terminates FL in advance, the accuracy loss will become marginal, as shown by Figures 17 and 18.
@@ 4.5.2 Independent implementation of the experiment methods
As a matter of fact, different kinds of efficient FL frameworks are compatible with each other and can be combined to further improve efficiency. For instance, FLrce can also transmit the compressed parameter updates to further reduce communication costs (early stopping + message compression), and Dropout can also reduce the local epochs to save computation resources (Dropout + accuracy relaxation), etc. To straightforwardly demonstrate the effectiveness of each method, the benchmarks are implemented independently in the experiment. We aim to find the optimal combination of the existing methods in future work.
@@ 4.5.3 Sub-model training does not necessarily reduce computation costs
A point worth mentioning is that Dropout does not effectively reduce the computation cost as Figure 11 shows. The possible reason is that Dropout mainly reduce the width (neurons) of a neural network rather than the depth (layers), while the training complexity mainly depends on the depth of the neural network. For example, if the model is trained with Pytorch, the computation cost mainly comes from the backward propagation process following the gradient graph that covers the entire neural network, while the size of the graph mainly depends on the depth of the neural network [45]. Therefore, Dropout does not effectively improve the computation efficiency. Comparatively, reducing the epoch/rounds of training is more helpful to reduce the computation cost, as shown by Figure 11, Fedprox and FLrce consume less energy than Dropout.
@@ 6 Conclusion
In this paper, we propose FLrce, an efficient FL framework excelling in overall communication and computation efficiency while maintaining accuracy with heterogeneous client contributions. Experiment results show that FLrce achieves desirable accuracy under unbalanced data distributions and significantly improves the efficiency of communication and computation resource utilization. In the future, we plan to investigate how to improve the efficiency for individual clients, e.g., by leveraging multi-task optimization [48], how to explore the relationships between clients more effectively, e.g., via learning vector quantization techniques [49], and conduct an in-depth theoretical study on FLrce.

### v4

@@ 4.2.1 Accuracy and Learning efficiency
4.2.2 Computation efficiency
We use energy consumption as a measurement of the computation cost [10, 14]. During the local training process, We measure the energy consumption of the Jetson Nano Developer Kits using a plug-in power monitor [44], when the training is completed, we sum up all clients’ energy consumption to derive the overall computation cost of the entire system, the results are available in Figure 11. As shown by Figure 11, FLrce consumes the least energy on all datasets, demonstrating the strength of FLrce in reducing the computation costs. Furthermore, motivated by [14], we combine accuracy and energy consumption to explore the computation efficiency of the benchmarks. Given the final accuracy and energy consumption, the computation efficiency is given by Equation (8):
As defined by Equation (8), the computation efficiency is proportional to accuracy and the inverse of energy consumption. It represents how a FL method is capable of utilizing the given computation resource to train an accurate global model. For a FL method, the higher accuracy it achieves and the less energy it consumes, the more computation efficient it will be. Figure 12 shows the computation efficiency of the benchmarks, for better visualization, the results have been normalized. As shown by Figure 12, FLrce obtains the highest computation efficiency on all of the four datasets. On average, FLrce improves the relative computation efficiency by at least 30%. This result demonstrates FLrce’s outstanding capability of achieving remarkable performance with limited computation resources.
@@ 4.2.2 Computation efficiency
4.2.3 Communication efficiency
From Figure 13, we can see that FLrce has the least bandwidth usage on EMNIST, Google Speech and CIFAR100. For CIFAR10, FLrce has the second least bandwidth usage that is slightly higher than Dropout. Even though Dropout uses less bandwidth than FLrce on CIFAR10, the final accuracy of Dropout is much lower than FLrce (14.9% vs. 57.8%). In this case, Dropout does not achieve a high communication efficiency as it does not maintain accuracy. Figure 14 shows that FLrce outperforms the baselines in terms of communication efficiency. On average, FLrce improves the relative communication efficiency by at least 43%.
4.3 Effect of the Early Stopping Threshold
We study the effect of the ES threshold $\psi$ empirically. We run FLrce multiple times with different values of $\psi$. As defined by subsection 3.3, to make sure that FLrce properly stops at a point with an acceptable validation accuracy, the ES mechanism should only be activated when a relatively large degree of conflict occurs among the selected clients, that is, the value of $\psi$ should not be too small. Furthermore, $\psi$ should not be too large as well; otherwise, the ES mechanism might not be activated in time and cannot reduce the resource consumption. Therefore, we tune the values of $\psi$ within $\{4.0,4.5,5.0,5.5,6.0,6.5,7.0\}$ using the well-known grid search method, and the results are shown by Figure 15, Figure 16 and Table IV. It is worth mentioning that the cases of $\psi<5$ and $\psi>6$ are not included in these results as they either terminate FLrce too early with low accuracy or fail to terminate before $T$. Moreover, Figure 15 and Figure 16 look exactly the same. This is because the computation/communication cost of FLrce per round is expected to be the same, so the computation/communication efficiency is proportional to $\frac{\text{accuracy}}{\text{round}}$ (see Equations (8) and (9)). Therefore, after normalization, the relative values of the computation and communication efficiency both become $\frac{\text{accuracy}}{\text{round}}$. For better illustration, we plot Figures 15 and 16 separately even though they show the same values. As shown by Table IV, FLrce successfully stops on all Datasets with $\psi=$ 5.0 or 5.5. For EMNIST and Google Speech, FLrce fails to stop before the maximum global iteration with $\psi\geq$ 6. Therefore, to make sure the ES mechanism be successfully activated, $\psi$ should be selected between $5\sim 5$ 5.5. As shown by Figure 15 and Figure 16, the value of $\psi$ to achieve the highest computation/communication efficiency is 5.0 (i.e. $\frac{P}{2}$, half of the number of selected clients per round). Accordingly, we set $\psi$ to 5.0 for EMNIST, CIFAR10, CIFAR100. For Google Speech, we set $\psi$ to 5.5 as it achieves higher accuracy than $\psi=5.0$ while still achieving the highest efficiency compared with the baselines (see Figures 12 and 14). Based on the empirical results, we suggest that resource-constrained IoT systems use a relatively small $\psi$ ($\approx$50of the active clients $P$) to maximize the efficiency of utilizing the limited resources. On the other hand, IoT systems with sufficient resource budget can use a relatively large threshold ($\approx$55$\sim$60of the active clients $P$) to achieve higher model accuracy. We aim to find the optimal value of $\psi$ via theoretical analysis in future works.
@@ 4.3 Effect of the Early Stopping Threshold
EMNIST 80.6% 44 82.2% 83 N/A N/A
Google Speech 60.5% 36 66.0% 52 N/A N/A
CIFAR10 57.8% 63 56.5% 81 56.4% 85
CIFAR100 37.1% 47 37.0% 48 36.9% 69
TABLE IV: Round and accuracy corresponding to different values of the Early Stopping threshold $\psi$. Note that values of $\psi$ that are too small ($<5$) or too large ($>6$) are not shown in this table. For one thing, if $\psi$ is too small, FLrce terminates in the very first few rounds with low final accuracy. For another thing, if $\psi$ is too large, the ES mechanism cannot be triggered before the maximum iteration $T$.
@@
Figure 15: Effect of the Early Stopping threshold on computation efficiency, the coordinates have been normalized for better visualization.
@@
Figure 16: Effect of the Early Stopping threshold on communication efficiency, the coordinates have been normalized for better visualization.
4.4 Effect of the Early Stopping Mechanism
To investigate whether the ES mechanism will cause undesirable accuracy loss, we compare the final accuracy of FLrce and FLrce without early stopping (FLrce w/o ES), using the threshold $\psi$ specified in the previous subsection. The only difference between FLrce and FLrce w/o ES is that the latter ignores the ES criterion and keeps training until reaching the maximum global iteration, and that relationship-based client selection strategy still applies in both settings. The comparison between FLrce and FLrce w/o ES is shown in Figure 10, Figure 17 and Figure 18. Note that Figures 17 and 18 present identical contents for the same reason as stated in the previous subsection. From Figures 17 and 18, we can see that FLrce can significantly increase the efficiency of the resource utilization with marginal accuracy sacrifice. For EMNIST and Google Speech, FLrce degrades the final accuracy by 2% and 4% respectively compared with FLrce w/o ES, while the efficiency of FLrce w/o ES is only about 45% and 55% of FLrce. Besides, compared with existing efficient FL frameworks, FLrce still maintains competitive performance even with the accuracy deduction, as shown by Figure 10 and Table III. For CIFAR10 and CIFAR100, FLrce even improves the final accuracy, as shown by Figures 17(c), 18(c), 17(d) and 18(d). One possible reason is that FLrce prevents the over-fitting of FL by terminating the FL process before the global model converges to some local sub-optimal solution.
@@ 4.5 Discussion
4.5.1 FLrce minimizes the accuracy loss caused by the trade-off between accuracy and efficiency
@@
For comparison, FLrce tackles this bottleneck from a novel perspective, while existing works reduce the resource consumption by reducing the computation/communication cost per round, FLrce reduces the overall resource consumption by reducing the total rounds of training. The advantage of this method is that FLrce does not have to make any trade-off (compress gradients, less epoch, etc.) in the training process, and can achieve a high accuracy faster given the same training rounds as the baselines. Consequently, when FLrce terminates FL in advance, the accuracy loss will become marginal, as shown by Figures 17 and 18.
@@ 4.5.2 Independent implementation of the experiment methods
As a matter of fact, different kinds of efficient FL frameworks are compatible with each other and can be combined to further improve efficiency. For instance, Dropout can further reduce the local epochs to save computation resources. FLrce can also transmit the compressed parameter updates to further reduce communication costs, or employ Dropout to overcome the computation/communication overheads of edge devices etc. To straightforwardly demonstrate the effectiveness of each method, the benchmarks are implemented independently in the experiment. We aim to find the optimal combination of the existing methods in future work.
@@ 4.5.3 Sub-model training does not necessarily reduce computation costs
A point worth mentioning is that Dropout does not effectively reduce the computation cost as Figure 11 shows. The possible reason is that Dropout mainly reduces the width (neurons) of a neural network rather than the depth (layers), while the training complexity mainly depends on the depth of the neural network. For example, if the model is trained with Pytorch, the computation cost mainly comes from the backward propagation process following the gradient graph that covers the entire neural network, while the size of the graph mainly depends on the depth of the neural network [46]. Therefore, Dropout does not effectively improve the computation efficiency. Comparatively, reducing the epoch/rounds of training is more helpful to reduce the computation cost, as shown by Figure 11, Fedprox and FLrce consume less energy than Dropout.
4.5.4 Generality of FLrce across various deep-learning tasks
In the experiment, FLrce exhibits the ability of maintaining computation and communication consistently across different types of deep-learning tasks, including image classification (EMNIST, CIFAR10, CIFAR100) and speech recognition (Google Speech). Firstly, as shown in Figure 10, the learning curve of FLrce on Google Speech presents similar patterns to that on EMNIST, CIFAR10 and CIFAR100. Secondly, as shown in Figures 12 and 14, FLrce achieves the highest computation and communication efficiency in both image classification and speech recognition tasks. Lastly, as shown in Table 3, the early stopping points where FLrce terminates prematurely in speech recognition (52) and image classification (44$\sim$63, averaging 51.3) are close. Overall, due to the generality displayed in the experiment, we expect to explore the potentiality of FLrce in broad areas of IoT applications in the future.
@@ 6 Conclusion
In this paper, we propose FLrce, an efficient FL framework excelling in overall communication and computation efficiency while maintaining accuracy with heterogeneous client contributions. Experiment results show that FLrce achieves remarkable accuracy under unbalanced data distributions and significantly improves the efficiency of communication and computation resources utilization.
The major limitation of FLrce is that it conducts full-model training on all clients regardless of the resource constraints of their edge devices. For devices that cannot train the entire global model because of limited capacity, this might become a bottleneck. As a solution, in the future, we aim to enhance the adaptability of FLrce to clients with limited device capacity. For example, consolidating FLrce with model pruning [25] and layer freezing [49] techniques to deal with resource-constrained devices’ incapability of training the entire global model.
In addition, we plan to investigate how to improve the efficiency for individual clients, e.g., by leveraging multi-task optimization [50], how to explore the relationships between clients more effectively, e.g., via learning vector quantization techniques [51], and conduct an in-depth theoretical study on FLrce.
Acknowledgement
This work is funded by the Australian Research Council under Grant No. DP220101823, DP200102611, and LP180100114.